OpenStudy (anonymous):

find the real or imaginary solutions by factoring x^4-10x^2+9=0

OpenStudy (jdoe0001):

\(\large x^4-10x^2+9=0\implies \begin{array}{cccllll} {\color{red}{ (x^2)}}^4&-10{\color{red}{ (x^2)}}&+9=0\\ &\uparrow &\uparrow \\ &-9-1&-9\cdot -1 \end{array}\)

OpenStudy (jdoe0001):

woops darn.. more typos =)

OpenStudy (jdoe0001):

\(\large x^4-10x^2+9=0\implies \begin{array}{cccllll} {\color{red}{ (x^2)}}^2&-10{\color{red}{ (x^2)}}&+9=0\\ &\uparrow &\uparrow \\ &-9-1&-9\cdot -1 \end{array}\)

OpenStudy (anonymous):

they exactly the same

OpenStudy (anonymous):

oh wait nvm

OpenStudy (anonymous):

but i just dont get like what you would do from there because i wasnt there the day the notes were taken and i just need something down on my paper for tomorrow

OpenStudy (jdoe0001):

there's an assumption you know how to factor a quadratic, thus

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

then those factors should factor it out :)

OpenStudy (jdoe0001):

then just solve for "x", is a quartic, thus you'd get 4 values for "x"